Each time that a certain alarm clock is set, it has a 95% chance of ringing properly. If it is set to ring every morning for five days, what is the probability that it rings properly every morning? it fails to ring at least one morning?
Answer by stanbon(75887)

Each time that a certain alarm clock is set, it has a 95% chance of ringing properly.
---------------------
Binomial Problem with n = 5 and p(proper) = 0.95 ; p(fails) = 0.05
--------
If it is set to ring every morning for five days, what is the probability that it rings properly every morning?
Ans:: P(x = 5) = 0.95^5 = 0.7738
------------------------------------
it fails to ring at least one morning?
Ans: P(fails at least once in 5) = 1- P(correct every day) = 0.2262
----------------
Cheers,
Stan H.
----------------
------------------
--------
P